A sample of pure PCl5 was introduced into an evacuated vessel at 473 K. After equilibrium was reached, the concentration of PCl5 was found to be 0.5 x 10-1 mol L-1. If Kc is 8.3 x 10-3 what are the concentrations of PCl3 and Cl2 at equilibrium?


1
Expert's answer
2020-12-06T10:34:58-0500

PCl5 = PCl3 + Cl2

c(PCl3) = c(Cl2) from stoichiometric coefficients

K = cPCl3*cCl2/cPCl5 = c2Cl2/cPCl5 = 8.3*10-3

cCl2 = cPCl3 = √(Kc*cPCl5) = √(8.3*10-3*0.5*10-1) = √(4.15*10-4) = 0.0204 mol/L